# Towards Graph Foundation Models: Learning Generalities Across Graphs via Task-Trees

### v2

@@ 4.2 GIT-S: Specification via Instruction Tuning
where $\psi$ is the instruction generation function for each task-tree, and $\kappa$ is the corresponding loss function. In this paper, as we use text-attributed graphs in our experiments, we define instructions as the embeddings of label descriptions encoded by a LLM, which is similar to Liu et al. (2024), and we use mean squared error as the loss function $\kappa$. Additional model analysis is provided in Appendix B.
@@ 5.1 Experimental Setup
Datasets. Our experiments are based on text-attributed graphs. Specifically, we include over 30 graphs spanning five domains: academic networks, e-commerce networks, knowledge graphs, molecular graphs, and temporal graphs. Detailed information is provided in Appendix E.1. During pretraining, we use citation network Arxiv, e-commerce network Products, knowledge graphs WN18RR and FB15K237, and molecular graphs Chemblpre and PCBA. During specialization, we use Arxiv for academic networks, Products for e-commerce networks, FB15K237 for knowledge graphs, and PCBA for molecular networks. For temporal graphs, which are e-commerce temporal graphs, we also use Products for SFT to assess robustness under temporal distribution shifts.
Baselines. We employ a wide range of baselines, including supervised GNNs like GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2019), self-supervised GNNs such as BGRL (Thakoor et al., 2022) and GraphMAE (Hou et al., 2022), graph foundation models like OFA (Liu et al., 2024). We also include some domain expert models. Details are provided in Appendix E.2.
Experimental Protocols. We use GraphSAGE (Hamilton et al., 2017) as the encoder, and each experiment is repeated at least five times with different seeds. We evaluate three settings: fine-tuning, in-context learning, and zero-shot learning. Fine-tuning is the standard approach where the model is fine-tuned on downstream tasks. In-context learning is a few-shot learning method without fine-tuning, where we randomly select $k$ samples from a given class, average these samples to form prototypes, and use the prototypes for classification. Following Liu et al. (2024); He & Hooi (2024), we randomly sample 500 5-way 3-shot tasks. If the number of classes is fewer than 5, the number of ways is set to the number of classes. Zero-shot learning operates similarly, but instead of using sample averages, it employs class description embeddings encoded by an LLM as the prototypes for prediction. Considering the metrics, we use accuracy in node classification and edge classification, and AUC in graph classification and link prediction. Additional details are provided in Appendix E.
@@ 5.2 Main Results
Domain-wise Results. Table 2 presents the experimental results averaged across each domain, with detailed results provided in Appendix F. The “held-out avg.” represents the average performance on all graphs, excluding those used in pretraining and specialization, ensuring an unbiased evaluation of downstream tasks. Notably, even the general GIT-G model surpasses baseline methods in certain settings. After specialization, the specialized GIT-S achieves additional performance gains, particularly in zero-shot and in-context learning. These findings align with our theoretical analysis, demonstrating that post-training a pretrained model enhances its adaptability to specialized domains.
Dataset-wise Results. Figure 3 illustrates the model performance across all graphs in the standard fine-tuning setting. We compare the performance of GIT against the strongest baseline methods. Notably, GIT consistently outperforms the best baselines across most datasets, further reinforcing the effectiveness of task-trees as learning instances.
Impact of Specialization. We summarize three key observations: (1) Specialization (GIT-S) enhances the performance of GIT-G across most settings (Table 2). However, the impact of specialization varies depending on the dataset used (Appendix Figure 9 and Table 23). (2) Specialization does not significantly degrade model performance on other domains (Table 25). (3) Applying the proposed specialization approach to other models also yields performance improvements in specific domains (Table 3), underscoring the potential of post-training in enhancing the capabilities of GNNs.
@@
5.3 Ablation on Training Strategies
@@
5.4 Comparison to Domain Experts
5.5 Comparison Between Task-Trees and Subgraphs
We evaluate the efficiency and performance of subgraphs and task-trees in the fine-tuning setting, as shown in Figure 4 and Table 4. To ensure a fair comparison, we implement a subgraph-based variant of GIT (GIT-SubG) by replacing task-trees with subgraphs. The results demonstrate that task-trees achieve both higher efficiency and improved model performance. While Table 4 reports results from three domains, the evaluation spans multiple tasks, including node classification (academic networks), edge classification (knowledge graphs), and graph classification (molecular networks).
6 Conclusion and Limitations
We introduce task-trees as basic learning instances for aligning graph-based tasks and provide both theoretical and empirical analyses to demonstrate their effectiveness. Building on task-trees, we develop GIT, a pretrained model that leverages a small set of graphs to enhance performance across more than 30 graphs spanning five domains. While our findings validate that task-trees preserve cross-task generalities at least partially, identifying the exact transferable patterns shared across graphs remains an open challenge. Further investigation from both empirical and theoretical perspectives is a promising direction for future research.
@@ Appendix A Related Work
Graph Neural Networks. GNNs are a class of learning models specifically designed to operate on graph-structured data and have demonstrated substantial success across a variety of domains. Their strength lies in their ability to perform relational learning, where information from neighboring nodes is aggregated and used to enhance node representations. For instance, GCN (Kipf & Welling, 2017) utilizes message-passing to aggregate information from neighboring nodes to central nodes. Building on this, models such as GraphSAGE (Hamilton et al., 2017) and GAT (Veličković et al., 2018) introduce innovative techniques like neighborhood sampling and attention mechanisms, respectively, further advancing performance on graph learning tasks. However, these methods are limited to solving a single task by training from the scratch.
@@
Graph Foundation Models. Foundation models are designed as general-purpose solvers capable of handling various tasks across different domains. For instance, LLMs, the foundation models in natural language processing, are capable of performing tasks such as summarization, translation, and entity recognition, as well as question-answering. However, building such versatile foundation models for graphs presents unique challenges due to the inherent feature, structural, and task heterogeneity across different graph domains and tasks. To address these challenges, Qiu et al. (2020) pretrained GNNs using subgraphs as basic units, mitigating structural heterogeneity. Building on this, Sun et al. (2023) reformulated node-, edge-, and graph-level tasks into subgraph-level tasks, tackling task heterogeneity. Additionally, Huang et al. (2023) and Liu et al. (2024) applied LLMs to unify the feature spaces of cross-domain graphs, addressing feature heterogeneity. These approaches enable models to operate on cross-domain and cross-task graphs. Further advancements, such as He & Hooi (2024) and Li et al. (2024), improve node embeddings by jointly optimizing GNN and LLM encoders, facilitating various downstream tasks like few-shot learning and zero-shot learning. However, most of these approaches rely on subgraphs as the primary learning instances, which can result in inefficient training and reduced expressiveness, as discussed in the main paper. Other efforts to resolve feature heterogeneity include methods like singular vector decomposition (SVD) (Zhao et al., 2024a; Yu et al., 2024), non-parametric encoders (Zhao et al., 2024b), or synthesizing text-attributed graphs (Wang et al., 2025). Notably, a recent study by Wang et al. (2024b) explores computation trees as transferable patterns in graphs. However, our work differs in three key aspects: (1) While Wang et al. (2024b) focus on identifying transferable patterns across graphs, our approach is centered on designing fundamental learning instances. (2) Our theoretical framework also serves as a foundational basis for Wang et al. (2024b). (3) Our proposed GIT significantly simplifies the model proposed by Wang et al. (2024b), demonstrating a minimally applicable design that retains effectiveness.
@@
Another line of research focuses on designing GFMs for single tasks or domains, thereby avoiding the complexities of feature, structural, or task heterogeneity. For example, Galkin et al. (2024) propose a foundation model for reasoning tasks on knowledge graphs, using triplets as basic transferable patterns. Zhao et al. (2023) introduce a foundation model for molecular graphs, employing LLMs to align semantics between datasets and encode key motifs. In node classification, Li et al. (2024) propose a zero-shot learning foundation model, while Zhao et al. (2024a) present a feature alignment method based on SVD for node-level graph foundation models. Recently, Zhao et al. (2024b) designed a foundation model for node classification using a non-parametric classifier. Meanwhile, Chen et al. (2024a), Tang et al. (2024), Guo et al. (2023), and Wang et al. (2024a) have explored using LLMs as graph reasoners to solve graph tasks, similar to their role in vision language models. While these methods excel at specific tasks or domains, they are not suitable as general graph solvers across diverse tasks. In contrast to these approaches, our proposed GIT model is pretrained on diverse task trees to acquire general reasoning capabilities, allowing it to quickly specialize in specific domains through instruction tuning.
@@ B.2 More Analysis on Domain Regularizer
How the Domain Regularizer Works? To better understand how the domain regularizer functions, we conduct an analysis to demonstrate its benefits in regulating domain distances while preserving task structures for each dataset. We use eight datasets provided by Liu et al. (2024) for pretraining: Cora, Pubmed, Arxiv, WikiCS, WN18RR, FB15K237, CHEMHIV, and CHEMPCBA. The analysis results are presented in Figure 5.
@@ B.5 Scaling Law
B.6 Limitations and Future Work.
We treat task-trees as basic learning instances in graphs, similar to how images are used in vision and text in language. Despite we have validated that cross-graph generalities are, at least partially, preserved in task-trees, identifying the exact transferable patterns shared across graphs remains a challenge. Further exploration of this issue from both empirical and theoretical perspectives is worthwhile. Additionally, for message-passing GNNs, converting graphs into tree structures inevitably leads to some loss of information, limiting the model expressiveness. Advanced GNNs (Morris et al., 2019, 2023) or techniques (Zhang et al., 2021) could be used to enhance expressiveness. Moreover, as GIT is a relatively simple model, investigating potential extensions is also a promising direction, as discussed in Appendix C.
@@ C.1 Pretraining
How to Design Reconstruction Tasks? Theorem 3.5 suggests that a well-designed encoder, capable of effectively handling reconstruction tasks during pretraining, can improve the model’s generalization ability. One approach is to use more powerful encoders to enhance reconstruction performance. Another approach is to introduce additional reconstruction losses to further refine the encoder. For example, methods such as those proposed by Qiu et al. (2020) and Hou et al. (2022), or designing more comprehensive reconstruction objectives could be explored.
@@ C.2 Specialization via Instruction Tuning
How to Leverage SFT Data? In scenarios with limited instructions, standard supervised fine-tuning may struggle to capture sufficient knowledge of the target domain. To address this, methods could be proposed to better utilize the unlabeled instances in the SFT dataset, thus enhancing model adaptation (Sohn et al., 2020).
@@ E.2 Baselines
OFA (Liu et al., 2024). A cross-task and cross-domain graph foundation model that treats subgraphs as the basic learning instances. It introduces a graph prompt learning framework to enable in-context and zero-shot learning.
@@ E.4 Evaluation Settings
In-context Learning. This is a kind of few-shot learning without fine-tuning the model parameters. We randomly select $k$ samples from a certain class, and average the selected samples to form prototypes, which is used for classification. We follow existing GFM works (Liu et al., 2024; He & Hooi, 2024) to conduct 500 randomly sampled 5-way 3-shot learning tasks. If the number of classes is less than 5, the number of ways is set to the number of classes.
@@ E.5 Hyper-Parameters
Baselines. For the baseline methods, we follow the hyperparameters reported in (Liu et al., 2024; Chen et al., 2024b). If the hyperparameters are not provided, we set the number of epochs to 1,000, the batch size to 4,096, early stopping at 200, and the hidden dimension to 768, using a 2-layer GraphSAGE as the backbone with batch normalization and ReLU activation. For optimization, we use AdamW with a weight decay of 1e-6 and tune the learning rate from 1e-3, 1e-4, 1e-5, reporting the best performance. For methods with attention mechanisms, we set 4 attention heads.
GIT. The model architecture and pretraining parameters of our GIT are presented in Table 7. The specific fine-tuning hyperparameters, categorized by domain, are shown in Tables 8, 9, 10, 11, and 12. For in-context learning and zero-shot learning results without fine-tuning, the general model does not involve any hyperparameters. For the specialized model, we tune the hyperparameters of SFT epochs from 10 to 500, in steps of 10, the SFT learning rate from 1e-4, 1e-5, 1e-6, 1e-7, 1e-8, and the normalization method from None, BN.

### v3

@@ 4.2 GIT-S: Specification via Instruction Tuning
where $\psi$ is the instruction generation function for each task-tree, and $\kappa$ is the corresponding loss function. In this paper, as we use text-attributed graphs in our experiments, we define instructions as the embeddings of label descriptions encoded by a LLM, which is similar to Liu et al. (2024a), and we use mean squared error as the loss function $\kappa$. Additional model analysis is provided in Appendix B.
@@ 5.1 Experimental Setup
Datasets. We conduct experiments on over 30 text-attributed graphs spanning five domains: academic networks, e-commerce networks, knowledge graphs, molecular graphs, and temporal graphs. Pretraining is performed on a diverse subset including Arxiv (academic), Products (e-commerce), WN18RR and FB15K237 (knowledge), and Chemblpre and PCBA (molecular). Specialization is evaluated on representative datasets for each domain: Arxiv, Products, FB15K237, and PCBA. For temporal graphs, which are e-commerce temporal graphs, we also use Products for SFT to assess robustness under temporal distribution shifts. We provide the full dataset details in Appendix E.1.
Baselines. We compare against a broad spectrum of baselines, including supervised GNNs (GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2019)), self-supervised models (BGRL (Thakoor et al., 2022), GraphMAE (Hou et al., 2022)), and graph foundation models (OFA (Liu et al., 2024a), GraphPrompt+ (Liu et al., 2023), All in One (Sun et al., 2023), OpenGraph (Xia et al., 2024), AnyGraph (Xia & Huang, 2024)). We also include domain-specific expert models for comparison. See Appendix E.2 for full descriptions.
Experimental Protocols. We use GraphSAGE (Hamilton et al., 2017) as the encoder and repeat each experiment five times with different random seeds. We evaluate three paradigms: fine-tuning, in-context learning, and zero-shot learning. Fine-tuning updates all model parameters on downstream data. In in-context learning (few-shot without fine-tuning), we randomly sample $k$ instances per class, compute prototype embeddings via averaging, and classify using nearest-prototype inference. Following Liu et al. (2024a); He & Hooi (2024), we sample 500 5-way 3-shot tasks; if the number of classes is fewer than 5, we use the actual number of classes as the number of ways. Zero-shot learning replaces prototypes with class description embeddings generated by an LLM. For evaluation, we report accuracy for node classification and edge classification, and AUC for graph classification and link prediction. Additional details are provided in Appendix E.
@@ 5.2 Main Results
Domain-Wise Performance. Table 2 reports average performance across five domains, with detailed per-dataset results provided in Appendix F. The “held-out avg.” denotes the average score on all graphs excluded from pretraining and specialization, offering an unbiased evaluation of generalization. Notably, the general-purpose model GIT-G already outperforms strong baselines in several domains. With domain-specific specialization, GIT-S achieves further gains—especially in zero-shot and in-context settings—highlighting the benefits of post-training in adapting pretrained models to specific domains. These results align with our theoretical findings, demonstrating enhanced adaptability through specialization.
Dataset-Wise Performance. Figure 3 shows fine-tuning performance across all datasets. GIT consistently outperforms the strongest baseline in the majority of cases, validating the effectiveness of task-trees as generalizable learning units across heterogeneous graph tasks.
Effect of Specialization. We summarize three key observations: (1) Specialization (GIT-S) enhances the performance of GIT-G across most settings (Table 2). However, the impact of specialization varies depending on the dataset used (Appendix Figure 9 and Table 25). (2) Specialization does not significantly degrade model performance on other domains (Table 27). (3) Applying the proposed specialization approach to other models also yields performance improvements in specific domains (Table 4), showing the generalization of post-training in enhancing model capacity.
5.3 Comparison to State-of-The-Art Methods
Table 3: Performance comparison to SOTA methods.
Academic KG Molecule
GraphPrompt+ 74.80 74.78 72.99
All in one 75.25 74.92 71.87
OpenGraph 74.64 71.38 72.84
AnyGraph 75.01 74.30 72.49
GIT - G 75.82 75.73 74.57
To evaluate the effectiveness of GIT, we compare it against recent graph foundation models, including GraphPrompt+ (Liu et al., 2023), All in One (Sun et al., 2023), OpenGraph (Xia et al., 2024), and AnyGraph (Xia & Huang, 2024), under the pretrain-then-finetune paradigm. As shown in Table 3, we benchmark GIT-G and all baselines across three representative domains: academic networks (node classification), knowledge graphs (edge classification), and molecular graphs (graph classification). GIT-G consistently outperforms all competing methods across tasks and domains. While many existing models pursue broad generalization by addressing multiple aspects—e.g., domain shifts, modality fusion, and prompt engineering—GIT adopts a targeted approach centered on task alignment. Its core innovation lies in the use of task-trees, which provide a unified abstraction across graph tasks.
5.4 Ablation on Training Strategies
@@ 5.5 Comparison to Domain Experts
Table 6: Results on non-text-attributed graphs.
Setting GraphMAE GIT-G
w/o SVD + w/o pretrain 95.02 94.27
w. SVD + w/o pretrain 94.89 95.50
w. SVD + w. pretrain 95.10 95.70
5.6 Task-Trees vs. Subgraphs: Efficiency and Performance
We compare the efficiency and effectiveness of task-trees against subgraph-based learning units in the fine-tuning setting, as shown in Figure 4 and Table 5. For a fair comparison, we implement a subgraph-based variant of our model, denoted GIT-SubG, by substituting task-trees with subgraphs while keeping all other components unchanged.
Experimental results across three domains—academic (node classification), knowledge graphs (edge classification), and molecular graphs (graph classification)—demonstrate that task-trees consistently outperform subgraphs in both computational efficiency (Figure 4) and predictive performance (Table 5). This highlights the advantage of task-trees in serving as compact, expressive, and structurally aligned learning instances, especially in cross-task and cross-domain generalization scenarios.
5.7 Results on Non-Text-Attributed Graphs
While prior experiments use text-attributed graphs, GIT does not inherently rely on textual information. We adopt text-attributed datasets to isolate the impact of task heterogeneity, while sidestepping the confounding effects of feature heterogeneity. Textual attributes enable feature alignment across graphs via a shared encoder, allowing us to more clearly assess the benefits of task-tree generalization. Importantly, GIT is compatible with non-textual graphs. To demonstrate this, we introduce a lightweight module that addresses feature heterogeneity by applying SVD to project features into a shared space. As shown in Table 6, we pretrain GIT-G on non-text-attributed datasets—PubMed (node classification), Citeseer (link prediction), and IMDB-B (graph classification)—and fine-tune on Cora (link prediction). Results, evaluated via AUC, confirm that GIT-G remains effective without relying on textual information, validating its broader applicability.
6 Conclusion and Limitations
@@
We introduce task-trees as unified learning instances for aligning heterogeneous graph-based tasks, supported by both theoretical and empirical analyses. Based on this abstraction, we develop GIT, a pretrained graph foundation model that leverages a small set of source graphs to generalize across over 30 datasets spanning five diverse domains. Our results demonstrate that task-trees preserve transferable patterns across tasks, offering a scalable and principled pathway for cross-domain generalization in graph learning.
Limitations and Future Work. While task-trees serve as a promising abstraction—analogous to images in vision and sentences in language—several challenges remain. First, although our results suggest that task-trees capture generalities across tasks, pinpointing the precise transferable patterns remains an open question. Future work should explore this from both theoretical and empirical perspectives. Second, transforming graph neighborhoods into tree structures may lead to information loss due to the limitations of message-passing GNNs. This can affect model expressiveness, especially for graphs with complex higher-order structures. Incorporating more expressive architectures, such as higher-order GNNs (Morris et al., 2019, 2023) or advanced graph modeling (Wang et al., 2025b), may help mitigate this issue. Finally, while GIT demonstrates strong performance with a lightweight design, exploring more sophisticated variants—e.g., incorporating attention mechanisms, graph prompting, or task-conditioned decoders—offers exciting directions for extending the generality and adaptability of graph foundation models. We outline several of these opportunities in Appendix C.
Acknowledgments
This work was partially supported by the NSF under grants IIS-2321504, IIS-2334193, IIS-2340346, IIS-2217239, CNS-2426514, and CMMI-2146076, and Notre Dame Strategic Framework Research Grant (2025). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Appendix A Related Work
Graph Neural Networks. GNNs are a class of learning models specifically designed to operate on graph-structured data and have demonstrated substantial success across a variety of domains. Their strength lies in their ability to perform relational learning, where information from neighboring nodes is aggregated and used to enhance node representations. For instance, GCN (Kipf & Welling, 2017) utilizes message-passing to aggregate information from neighboring nodes to central nodes. Building on this, models such as GraphSAGE (Hamilton et al., 2017) and GAT (Veličković et al., 2018) introduce innovative techniques like neighborhood sampling and attention mechanisms, respectively, further advancing performance on graph learning tasks. However, these methods are limited to solving a single task by training from the scratch (Zhang et al., 2019; Fan et al., 2022; Qian et al., 2025; Ma et al., 2023; Qian et al., 2024).
@@
Graph Foundation Models. Foundation models are designed as general-purpose solvers capable of handling various tasks across different domains. For instance, LLMs, the foundation models in natural language processing, are capable of performing tasks such as summarization, translation, and entity recognition, as well as question-answering. However, building such versatile foundation models for graphs presents unique challenges due to the inherent feature, structural, and task heterogeneity across different graph domains and tasks. To address these challenges, Qiu et al. (2020) pretrained GNNs using subgraphs as basic units, mitigating structural heterogeneity. Building on this, Sun et al. (2023) reformulated node-, edge-, and graph-level tasks into subgraph-level tasks, tackling task heterogeneity. Additionally, Huang et al. (2023) and Liu et al. (2024a) applied LLMs to unify the feature spaces of cross-domain graphs, addressing feature heterogeneity. These approaches enable models to operate on cross-domain and cross-task graphs. Further advancements, such as He & Hooi (2024) and Li et al. (2024), improve node embeddings by jointly optimizing GNN and LLM encoders, facilitating various downstream tasks like few-shot learning and zero-shot learning. However, most of these approaches rely on subgraphs as the primary learning instances, which can result in inefficient training and reduced expressiveness, as discussed in the main paper. Other efforts to resolve feature heterogeneity include methods like singular vector decomposition (SVD) (Zhao et al., 2024a; Yu et al., 2024), non-parametric encoders (Zhao et al., 2024b), or synthesizing text-attributed graphs (Wang et al., 2025a). Notably, a recent study by Wang et al. (2024b) explores computation trees as transferable patterns in graphs. However, our work differs in three key aspects: (1) While Wang et al. (2024b) focus on identifying transferable patterns across graphs, our approach is centered on designing fundamental learning instances. (2) Our theoretical framework also serves as a foundational basis for Wang et al. (2024b). (3) Our proposed GIT significantly simplifies the model proposed by Wang et al. (2024b), demonstrating a minimally applicable design that retains effectiveness.
@@
Another line of research focuses on designing GFMs for single tasks or domains, thereby avoiding the complexities of feature, structural, or task heterogeneity. For example, Galkin et al. (2024) propose a foundation model for reasoning tasks on knowledge graphs, using triplets as basic transferable patterns. Zhao et al. (2023) introduce a foundation model for molecular graphs, employing LLMs to align semantics between datasets and encode key motifs. In node classification, Li et al. (2024) propose a zero-shot learning foundation model, while Zhao et al. (2024a) present a feature alignment method based on SVD for node-level graph foundation models. Recently, Zhao et al. (2024b) designed a foundation model for node classification using a non-parametric classifier. Meanwhile, Chen et al. (2024a), Tang et al. (2024), Guo et al. (2023), Liu et al. (2024b), and Wang et al. (2024a) have explored using LLMs as graph reasoners to solve graph tasks, similar to their role in vision language models. While these methods excel at specific tasks or domains, they are not suitable as general graph solvers across diverse tasks. In contrast to these approaches, our proposed GIT model is pretrained on diverse task trees to acquire general reasoning capabilities, allowing it to quickly specialize in specific domains through instruction tuning.
@@ B.2 More Analysis on Domain Regularizer
How the Domain Regularizer Works? To better understand how the domain regularizer functions, we conduct an analysis to demonstrate its benefits in regulating domain distances while preserving task structures for each dataset. We use eight datasets provided by Liu et al. (2024a) for pretraining: Cora, Pubmed, Arxiv, WikiCS, WN18RR, FB15K237, CHEMHIV, and CHEMPCBA. The analysis results are presented in Figure 5.
@@ C.1 Pretraining
How to Design Reconstruction Tasks? Theorem 3.5 suggests that a well-designed encoder, capable of effectively handling reconstruction tasks during pretraining, can improve the model’s generalization ability. One approach is to use more powerful encoders to enhance reconstruction performance. Another approach is to introduce additional reconstruction losses to further refine the encoder. For example, methods such as those proposed by Qiu et al. (2020), and Hou et al. (2022), Ju et al. (2023b), and Wen et al. (2024), or designing more comprehensive reconstruction objectives could be explored.
@@ C.2 Specialization via Instruction Tuning
How to Leverage SFT Data? In scenarios with limited instructions, standard supervised fine-tuning may struggle to capture sufficient knowledge of the target domain. To address this, methods could be proposed to better utilize the unlabeled instances in the SFT dataset, thus enhancing model adaptation (Sohn et al., 2020; Ni et al., 2025).
@@ E.2 Baselines
OFA (Liu et al., 2024a). A cross-task and cross-domain graph foundation model that treats subgraphs as the basic learning instances. It introduces a graph prompt learning framework to enable in-context and zero-shot learning.
@@ E.4 Evaluation Settings
In-context Learning. This is a kind of few-shot learning without fine-tuning the model parameters. We randomly select $k$ samples from a certain class, and average the selected samples to form prototypes, which is used for classification. We follow existing GFM works (Liu et al., 2024a; He & Hooi, 2024) to conduct 500 randomly sampled 5-way 3-shot learning tasks. If the number of classes is less than 5, the number of ways is set to the number of classes.
@@ E.5 Hyper-Parameters
Baselines. For the baseline methods, we follow the hyperparameters reported in (Liu et al., 2024a; Chen et al., 2024b). If the hyperparameters are not provided, we set the number of epochs to 1,000, the batch size to 4,096, early stopping at 200, and the hidden dimension to 768, using a 2-layer GraphSAGE as the backbone with batch normalization and ReLU activation. For optimization, we use AdamW with a weight decay of 1e-6 and tune the learning rate from 1e-3, 1e-4, 1e-5, reporting the best performance. For methods with attention mechanisms, we set 4 attention heads.
GIT. The model architecture and pretraining parameters of our GIT are presented in Table 9. The specific fine-tuning hyperparameters, categorized by domain, are shown in Tables 10, 11, 12, 13, and 14. For in-context learning and zero-shot learning results without fine-tuning, the general model does not involve any hyperparameters. For the specialized model, we tune the hyperparameters of SFT epochs from 10 to 500, in steps of 10, the SFT learning rate from 1e-4, 1e-5, 1e-6, 1e-7, 1e-8, and the normalization method from None, BN.